OpenStudy (anonymous):

Solve: y/-2 + 5 = 13

OpenStudy (anonymous):

I worked out the problem but unsure if I'm even close to the right answer. Is it -16?

OpenStudy (phi):

is the problem \[ \frac{y}{-2} + 5 = 13\] ?

OpenStudy (anonymous):

Yes it is, sorry should have clarified that. It's a two-step equation.

OpenStudy (phi):

you can check your answer. replace y with -16 and see if the equation is true. \[\frac{-16}{-2} + 5 = 13 ? \] -16/-2 is +8 so you have \[ 8+5 = 13 ? \] 8+5 is 13 so you get \[ 13 = 13 \] which is true. so y = -16 is correct

OpenStudy (anonymous):

Thanks so much @phi also for future knowledge how do you write the fraction?

OpenStudy (anonymous):

alt code?

OpenStudy (phi):

to solve \[ \frac{y}{-2} + 5 = 13 \] add -5 to both sides \[ \frac{y}{-2} + 5 + -5 = 13 +- 5\] simplify to get \[ \frac{y}{-2} = 8 \] multiply both sides by -2 to get y = -8 for fractions, use the equation editor (bottom left button)

OpenStudy (phi):

**multiply both sides by -2 to get y = -16

OpenStudy (anonymous):

Thanks a bunch! :D